OpenStudy (larrylovestein):

(4/9)^-1= A) -9/4 B) -4/9 C) 9/4

OpenStudy (anonymous):

1/( 4/9 ) = 9/4 Invert the fraction and then multiply 1*9/4 = 9/4 or C.

OpenStudy (larrylovestein):

thanks